OpenStudy (anonymous):

How do you find out how many square feet a bag of concrete will cover if you want a slab that is 6inches thick and concrete is sold in cubed yards?

OpenStudy (phi):

change cubic yards to cubic feet 3 ft / yard cube everything: (3 ft)^3/yard^3 or 27 ft^3 per yard^2 change 6 inches to feet: 0.5 feet (1/2 of a foot) volume is length*width*height or L*W*H L*W is area (in square feet), so that is what you are looking for. call L*W= A (for area) so A*H is the volume. H= 0.5 ft A*0.5= 27 ft^3 A= 54 ft^2

OpenStudy (anonymous):

thanks!